OpenStudy (anonymous):

In a chemistry class, 6 liters of a 4% silver iodide must be mixed with a 10% solution to get a 6% solution. How many liters of the 10% are needed?

OpenStudy (anonymous):

.04(6L)+.1(x)=.06(6+x) 4(6)+10x=6(6+x) 24+10x=36+6x 4x=12 x=3 liters

OpenStudy (anonymous):

ur quick. i just finished he is right

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

you are great!!!

OpenStudy (anonymous):

what he did was compare the percentage of acid 04(6L)+.1x is the amount of acid so if we divide that by the amount of solution: 6+x we get the percentage which should b equal to 6% so: (.04(6)+.1x)/(6+x)=.06

OpenStudy (anonymous):

You make it sound so easy, thanks

OpenStudy (anonymous):

lol it takes practice